OpenStudy (anonymous):

Write an equation in slope-intercept form of the line with slope of -2 that contains (3, -7)

OpenStudy (shinalcantara):

Slope-intercept form: \[y = mx + b\] ------------------ Using point-slope form (y-y1)=m(x-x1) x1=3 y1=-7 m=-2 ------------------- (y-y1)=m(x-x1) y-(-7)=-2(x-3) y+7 = -2x + 6 y = -2x + 6 - 7 y = -2x - 1